Find the value of r so the line that passes through each pair of points has the given slope.
(r,5), ( -2, r) ,m = -2/9 is a fraction )))


Answer by stanbon(75887)

Find the value of r so the line that passes through each pair of points has the given slope.
(r,5), ( -2, r) ,m = -2/9 is a fraction
===========================
slope = (r-5)/(-2-r) = -2/9
---
Cross-multiply:
9(r-5) = 2(2+r)
---
9r-45 = 4+2r
7r = 49
---
r = 7
=========================
Cheers,
Stan H.
====================